# On the speed of coming down from infinity for (sub)critical branching processes with pairwise interactions

### v1

@@ 3.4 Proof of Theorem 3
It only remains to prove (63). Consider the function $F:(-1,\infty)\rightarrow\mathbb{R}$ given by $F(x)=\log(1+x)$. The function $F$ can be extended in such a way that $F\in C^{2}(\mathbb{R},\mathbb{R})$. Indeed, we may consider $\varphi F$, where $\varphi\in C^{\infty}(\mathbb{R},\mathbb{R})$ is equal to 00 on $(-\infty,-1]$ and to $1$ on $\overline{D\cup\mathbb{N}_{0}}$, thanks to the smooth Urysohn lemma since the aforementioned sets are disjoint and closed. In other words, we may extend F from $\overline{D\cup\mathbb{N}_{0}}$ to $\mathbb{R}$ such that $F\in C^{2}(\mathbb{R},\mathbb{R})$. Then, by applying Itô’s formula ([18, Chapter 2, Theorem 5.1]) and Proposition 1, for $t\geq 0$,
$\log(1+Z_{t}^{(z)})=\log(1+z)+\int_{0}^{t}\int_{\mathcal{X}}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$
Next, let us consider the functions $h$ and $\psi$ as in Lemma 2. Then, the process $(\log(1+Z_{t}^{(z)}))_{t\geq 0}$ is a semimartingale that takes values in $\{\log(1),\log(2),\log(3),\dots\}$ and can be written as
$\log(1+Z_{t}^{(z)})=\log(1+z)+\int_{0}^{t}\hat{\psi}(\log(1+Z_{s}^{(z)})){\rm d}s+A_{t}+M^{\rm d}_{t},\quad t\geq 0,$
where $\hat{\psi}(y):=\psi(y)\mathbf{1}_{\{y\in\log(1+D)\}}$, for $y\in\mathbb{R}$ and $\log(1+D):=\{\log(1+x):x\in D\}$, is a Borel locally bounded function by Lemma 2,
$A_{t}:=\int_{0}^{t}h(Z_{s}^{(z)})\mathbf{1}_{\{\log(1+Z_{s}^{(z)})\not\in\log(1+D)\}}{\rm d}s$
@@
$M^{\rm d}_{t}:=\int_{0}^{t}\int_{\mathcal{X}}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)$
is a càdlàg local martingale. Indeed, by (58) in Lemma 12, $(M^{\rm d}_{t})_{t\geq 0}$ is a locally square-integrable martingale.
@@
$y_{0}=\log(1+z),\quad y_{t}^{\prime}=\hat{\psi}(y_{t})=\psi(y_{t}),$
since $\psi=\hat{\psi}$ on $\log(1+D)$. This flow is thus associated with the vector field $\hat{\psi}$ which is locally Lipschitz on $\log(1+D)$ by Lemma 2. Set $T^{\prime}(y_{0})=T(z)$, and for $\varepsilon>0$, let
$T_{\log(1+D),\varepsilon}(y_{0})\coloneqq\sup\left\{t\in[0,T^{\prime}(y_{0})):\,\forall s\leq t,\,\bar{B}(y_{s},\varepsilon)\cap\log(1+\mathbb{N}_{0})\subset\log(1+D)\right\}$
be the last time when $y_{t}$ starting from $y_{0}$ is at distance $\varepsilon>0$ from the boundary of $\log(1+D)$ under the Euclidean distance. Note that $T_{\log(1+D),\varepsilon}(y_{0})\geq T_{D,\varepsilon}(z)$ and that $\hat{\psi}$ is $(0,\alpha)$ nonexpansive on $\log(1+D)$ by Lemma 2. By applying [3, Lemma 2.1] to $(\log(1+Z_{t}^{(z)}))_{t\geq 0}$ with $\log(1+Z_{0}^{(z)})=y_{0}=\log(1+z)$ and the union bound, we obtain that, for any $t<T_{D,\varepsilon}(z)\wedge\varepsilon/4\kappa\wedge\eta_{\theta}$,
@@
$\quad\quad+\mathbb{P}\left(\sum_{u\leq t}\mathbf{1}_{\{S_{u-}\leq\varepsilon\}}|\Delta\log(1+Z_{u}^{(z)})|^{2}\geq\frac{\varepsilon^{2}}{4\sqrt{2}}\right),$ (65)
@@
$\mathbb{P}\left(\sup_{u\leq t}\int_{0}^{u}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}(\log(1+Z_{s-}^{(z)})-y_{s}){\rm d}A_{s}\geq\frac{\varepsilon^{2}}{16\sqrt{2}}\right)=0.$ (66)
@@
$Y_{r}^{(\varepsilon)}:=\int_{0}^{r}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}(\log(1+Z_{s-}^{(z)}))-y_{s}){\rm d}M^{\rm d}_{s},\quad\text{for}\quad r\geq 0.$
By (58) in Lemma 12, the process $(Y_{r}^{(\varepsilon)})_{r\geq 0}$ is a locally square-integrable martingale. On the other hand, for $p\geq 2$, we note by (58) in Lemma 12 that the (predictable) compensator $(V^{(p)}_{r})_{r\geq 0}$ of the process $(\sum_{s\leq r}|\Delta Y_{s}^{(\varepsilon)}|^{p})_{r\geq 0}$ satisfies $V^{(p)}_{r}\leq\frac{p!}{2}(\frac{\varepsilon}{\ln\theta})^{p}C_{\theta}r$, for $r\geq 0$, where
$C_{\theta}=2\left(d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i}\right).$ (67)
Then Lemma 22 implies that, for any $t<T_{D,\varepsilon}(z)\wedge\varepsilon/4\kappa\wedge\eta_{\theta}$,
@@
$\leq\exp\left(-\frac{\varepsilon\ln\theta}{64\sqrt{2}}\right),$ (68)
@@
$\sum_{s\leq r}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}|\Delta\log(1+Z_{s}^{(z)}))|^{2}=\int_{0}^{r}\int_{\mathcal{X}}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}\Big{(}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\Big{)}^{2}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x),$
for $r\geq 0$. In particular, by (58) in Lemma 12, the process $(\hat{Y}_{r}^{(\varepsilon)})_{r\geq 0}$, given by
$\hat{Y}_{r}^{(\varepsilon)}:=\int_{0}^{r}\int_{\mathcal{X}}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}\Big{(}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\Big{)}^{2}\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x),$
is a locally square-integrable martingale. By inequality (58) in Lemma 12, we see
$\int_{0}^{r}\int_{\mathcal{X}}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}\Big{(}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\Big{)}^{2}q({\rm d}u,{\rm d}y,{\rm d}x){\rm d}s\leq\frac{C_{\theta}}{(\ln\theta)^{2}}r,$ (69)
for $r\geq 0$. For $p\geq 2$, inequality (59) in Lemma 12 implies that the (predictable) compensator $(\hat{V}^{(p)}_{r})_{r\geq 0}$ of $(\sum_{s\leq r}|\Delta\hat{Y}_{s}^{(\varepsilon)}|^{p})_{t\geq 0}$ satisfies $\hat{V}^{(p)}_{r}\leq\frac{p!}{2}(\frac{2}{\ln\theta})^{p}C_{\theta}r$, for $r\geq 0$, where $C_{\theta}$ is as in (67). Then, it follows from (69) and Lemma 22 that, for $t<T_{D,\varepsilon}(z)\wedge\varepsilon/4\kappa\wedge\eta_{\theta}$,
@@
$\leq\exp\left(-\frac{\varepsilon\ln\theta}{64\sqrt{2}}\right),$ (70)
@@
Finally, (63) follows by combining (3.4), (66), (3.4) and (3.4). ∎
We continue with the proof of Theorem 3.
Proof of Theorem 3.
@@
$\lim_{t\downarrow 0}\mathbb{E}\left[\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}-1\right|^{p}\right]=0.$ (71)
@@
$-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1=-\mathbf{m}_{\rm int}s(v_{s}+1)\Big{(}\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}-1\Big{)}-\mathbf{m}_{\rm int}s(v_{s}+1)+s\mathbf{m}_{\rm int}-1$
and thus, (71) and Lemma 4 imply our claim in Theorem 3.
Now we prove (71). Note that
$\sup_{s\leq t}\left|\log\left(\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right)\right|=\log\left(\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right|\vee\sup_{s\leq t}\left|\frac{v_{s}+1}{Z_{s}^{(\infty)}+1}\right|\right),\quad t\geq 0.$
Fix $0\leq t<T(\infty)$, and choose $\varepsilon>e$ (i.e., $\ln\varepsilon>1$) such that $t<T(\infty)\wedge(\ln\varepsilon)/4\kappa\wedge\eta_{\theta}$. It follows from Lemma 13 that, for $u>\varepsilon$,
$\mathbb{P}\left(\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right|\geq u\right)\leq e^{-\frac{(\ln u)\ln\theta}{64\sqrt{2}}}=u^{-\frac{\log\theta}{64\sqrt{2}}};$
to see this note that $(\ln\varepsilon)/4\kappa\leq(\ln u)/4\kappa$, and thus $t<T(\infty)\wedge(\ln u)/4\kappa\wedge\eta_{\theta}$. Since $\theta>e^{128\sqrt{2}}$ by (${\bf H4}$), then, for $p\in[1,2]$, there is a constant $0<C_{1}(p)<\infty$ such that
$\mathbb{E}\left[\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right|^{p}\right]<C_{1}(p).$ (72)
@@
$\mathbb{E}\left[\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}-1\right|^{p}\right]<C_{2}(p).$
Therefore, for $p\in[1,2]$, (71) follows by Lemma 4, (42) and the dominated convergence theorem. The case $p\in(0,1)$ follows by Jensen’s inequality. This completes the proof. ∎
Corollary 2.
@@
    $\mathbb{E}\Big{[}\sup_{s\leq t}\Big{|}-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}\Big{|}^{p}\Big{]}<C(p),\quad\text{for}\quad t\leq T.$
@@
    $\mathbb{E}\Big{[}\sup_{r\leq s\leq t}Z_{s}^{(\infty)}\Big{]}<\infty\qquad\text{and}\qquad\mathbb{E}\Big{[}\sup_{r\leq s\leq t}(Z_{s}^{(\infty)})^{2}\Big{]}<\infty.$
@@
$-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}\leq-\mathbf{m}_{\rm int}s(v_{s}+1)\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}.$ (73)
Recall that $-\mathbf{m}_{\rm int}>0$. By Lemma 4, $\lim_{s\downarrow 0}-\mathbf{m}_{\rm int}s(v_{s}+1)=1$. In particular, $s\mapsto-\mathbf{m}_{\rm int}s(v_{s}+1)$ is bounded on $[0,T]$. Then, (73), Lemma 4 and (72) imply (i) for $p\in[1,2]$. For $p\in(0,1)$, the inequality in (i) follows by Jensen’s inequality. To prove (ii), note that, for $0<r\leq s\leq t$,
$Z_{s}^{(\infty)}\leq(-r\mathbf{m}_{\rm int})^{-1}\sup_{r\leq u\leq t}\left|-\mathbf{m}_{\rm int}uZ_{u}^{(\infty)}\right|.$
@@
3.5 Proof of Theorem 4
In this section, we prove Theorem 4. We adopt the methodology outlined in [29] (see also [30]), with modifications to accommodate the branching and cooperative terms of the BPI process. Recall the definition of $T(\infty)$ given before (33).
Lemma 14.
@@ 3.5 Proof of Theorem 4
$Z_{t}^{(\infty)}=Z_{r}^{(\infty)}+\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,x)N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$
@@
Recall that (${\bf H4}$) implies $\mathbf{m}_{\rm br}<\infty$, the latter together with $\mathbf{m}_{\rm int}<0$ implies that the BPI process does not explode in finite time; see [7, Example 1]. By Proposition 1, for all $z\in\mathbb{N}_{0}$ and $0<r\leq t<T(\infty)$,
@@
$\quad\quad\quad+\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(z)},u,y,x)\mathbf{1}_{\{x\geq 0\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (74)
Proposition 2 and Theorem 1 imply that, for any $0<t<T(\infty)$, $\lim_{z\rightarrow\infty}Z_{t}^{(z)}=Z_{t}^{(\infty)}$ (increasingly), almost surely. Recall that the Poisson random measure $N$ is an almost-surely $\sigma$-finite measure. Since $z\in\mathbb{R}_{+}\mapsto H(z,u,y,x)\mathbf{1}_{\{x\geq 0\}}$ is an increasing function, the monotone convergence theorem implies that
$\lim_{z\rightarrow\infty}\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(z)},u,y,x)\mathbf{1}_{\{x\geq 0\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)=\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,x)\mathbf{1}_{\{x\geq 0\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$
On the other hand, $|H(Z_{s-}^{(z)},u,y,-1)|\mathbf{1}_{\{x=-1\}}\leq|H(Z_{s-}^{(\infty)},u,y,-1)|\mathbf{1}_{\{x=-1\}}$, for all $r\leq s\leq t<T(\infty)$ and $(u,y,x)\in\mathcal{X}$. By Corollary 2 (ii), one can easily deduce that $|H(Z_{s-}^{(\infty)},u,y,-1)|\mathbf{1}_{\{x=-1\}}$ is integrable with respect to the Poisson random measure $N$, almost surely. Therefore, by the dominated convergence theorem,
$\lim_{z\rightarrow\infty}\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(z)},u,y,-1)\mathbf{1}_{\{x=-1\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)=\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,-1)\mathbf{1}_{\{x=-1\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$
@@
Let $\tilde{N}$ be the compensated Poisson point measure $N$ on +×$\mathbb{R}_{+}\times\mathcal{X}$ defined in (16), where $\mathcal{X}\coloneqq\{1,2\}\times(\mathbb{N}\cup\Delta)\times\mathbb{N}_{-1}$.
Lemma 15.

### v2

@@ 3.1 Integral representation of BPI processes
It only remains to prove (3.43). Consider the function $F:(-1,\infty)\rightarrow\mathbb{R}$ given by $F(x)=\log(1+x)$. The function $F$ can be extended in such a way that $F\in C^{2}(\mathbb{R},\mathbb{R})$. Indeed, we may consider $\varphi F$, where $\varphi\in C^{\infty}(\mathbb{R},\mathbb{R})$ is equal to 0 on $(-\infty,-1]$ and to $1$ on $\overline{D\cup\mathbb{N}_{0}}$, thanks to the smooth Urysohn lemma since the aforementioned sets are disjoint and closed. In other words, we may extend F from $\overline{D\cup\mathbb{N}_{0}}$ to $\mathbb{R}$ such that $F\in C^{2}(\mathbb{R},\mathbb{R})$. Then, by applying Itô’s formula (Theorem 5.1 in Chapter 2 of [17]) and Proposition 3.1, for $t\geq 0$,
$\log(1+Z_{t}^{(z)})=\log(1+z)+\int_{0}^{t}\int_{\mathcal{X}}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.47)
Next, let us consider the functions $h$ and $\psi$ as in Lemma 3.4. Then, the process $(\log(1+Z_{t}^{(z)}))_{t\geq 0}$ is a semimartingale that takes values in $\{\log(1),\log(2),\log(3),\dots\}$ and can be written as
$\log(1+Z_{t}^{(z)})=\log(1+z)+\int_{0}^{t}\hat{\psi}(\log(1+Z_{s}^{(z)})){\rm d}s+A_{t}+M^{\rm d}_{t},\quad t\geq 0,$ (3.48)
where $\hat{\psi}(y):=\psi(y)\mathbf{1}_{\{y\in\log(1+D)\}}$, for $y\in\mathbb{R}$ and $\log(1+D):=\{\log(1+x):x\in D\}$, is a Borel locally bounded function by Lemma 3.4,
$A_{t}:=\int_{0}^{t}h(Z_{s}^{(z)})\mathbf{1}_{\{\log(1+Z_{s}^{(z)})\not\in\log(1+D)\}}{\rm d}s$ (3.49)
@@
$M^{\rm d}_{t}:=\int_{0}^{t}\int_{\mathcal{X}}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)$ (3.50)
is a càdlàg local martingale. Indeed, by (3.35) in Lemma 3.9, $(M^{\rm d}_{t})_{t\geq 0}$ is a locally square-integrable martingale.
@@
$y_{0}=\log(1+z),\quad y_{t}^{\prime}=\hat{\psi}(y_{t})=\psi(y_{t}),$ (3.51)
since $\psi=\hat{\psi}$ on $\log(1+D)$. This flow is thus associated with the vector field $\hat{\psi}$ which is locally Lipschitz on $\log(1+D)$ by Lemma 3.4. Set $T^{\prime}(y_{0})=T(z)$, and for $\varepsilon>0$, let
$T_{\log(1+D),\varepsilon}(y_{0})\coloneqq\sup\left\{t\in[0,T^{\prime}(y_{0})):\,\forall s\leq t,\,\bar{B}(y_{s},\varepsilon)\cap\log(1+\mathbb{N}_{0})\subset\log(1+D)\right\}$ (3.52)
be the last time when $y_{t}$ starting from $y_{0}$ is at distance $\varepsilon>0$ from the boundary of $\log(1+D)$ under the Euclidean distance. Note that $T_{\log(1+D),\varepsilon}(y_{0})\geq T_{D,\varepsilon}(z)$ and that $\hat{\psi}$ is $(0,\alpha)$ nonexpansive on $\log(1+D)$ by Lemma 3.4. By applying Lemma 2.1 in [3] to $(\log(1+Z_{t}^{(z)}))_{t\geq 0}$ with $\log(1+Z_{0}^{(z)})=y_{0}=\log(1+z)$ and the union bound, we obtain that, for any $t<T_{D,\varepsilon}(z)\wedge\varepsilon/4\kappa\wedge\eta_{\theta}$,
@@
$\quad\quad+\mathbb{P}\left(\sum_{u\leq t}\mathbf{1}_{\{S_{u-}\leq\varepsilon\}}|\Delta\log(1+Z_{u}^{(z)})|^{2}\geq\frac{\varepsilon^{2}}{4\sqrt{2}}\right),$ (3.53)
@@
$\mathbb{P}\left(\sup_{u\leq t}\int_{0}^{u}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}(\log(1+Z_{s-}^{(z)})-y_{s}){\rm d}A_{s}\geq\frac{\varepsilon^{2}}{16\sqrt{2}}\right)=0.$ (3.54)
@@
$Y_{r}^{(\varepsilon)}:=\int_{0}^{r}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}(\log(1+Z_{s-}^{(z)}))-y_{s}){\rm d}M^{\rm d}_{s},\quad\text{for}\quad r\geq 0.$ (3.55)
By (3.35) in Lemma 3.9, the process $(Y_{r}^{(\varepsilon)})_{r\geq 0}$ is a locally square-integrable martingale. On the other hand, for $p\geq 2$, we note by (3.35) in Lemma 3.9 that the (predictable) compensator $(V^{(p)}_{r})_{r\geq 0}$ of the process $(\sum_{s\leq r}|\Delta Y_{s}^{(\varepsilon)}|^{p})_{r\geq 0}$ satisfies $V^{(p)}_{r}\leq\frac{p!}{2}(\frac{\varepsilon}{\ln\theta})^{p}C_{\theta}r$, for $r\geq 0$, where
$C_{\theta}=2\left(d\theta+c\theta+\sum_{i\geq 1}\theta^{i}\pi_{i}+\sum_{i\geq 1}\theta^{i}b_{i}\right).$ (3.56)
Then Lemma A.1 implies that, for any $t<T_{D,\varepsilon}(z)\wedge\varepsilon/4\kappa\wedge\eta_{\theta}$,
@@
$\leq\exp\left(-\frac{\varepsilon\ln\theta}{64\sqrt{2}}\right),$ (3.57)
@@
$\sum_{s\leq r}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}|\Delta\log(1+Z_{s}^{(z)}))|^{2}=\int_{0}^{r}\int_{\mathcal{X}}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}\Big{(}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\Big{)}^{2}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x),$ (3.58)
for $r\geq 0$. In particular, by (3.35) in Lemma 3.9, the process $(\hat{Y}_{r}^{(\varepsilon)})_{r\geq 0}$, given by
$\hat{Y}_{r}^{(\varepsilon)}:=\int_{0}^{r}\int_{\mathcal{X}}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}\Big{(}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\Big{)}^{2}\tilde{N}({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x),$ (3.59)
is a locally square-integrable martingale. By inequality (3.35) in Lemma 3.9, we see
$\int_{0}^{r}\int_{\mathcal{X}}\mathbf{1}_{\{S_{s-}\leq\varepsilon\}}\Big{(}\log\Big{(}1+\frac{H(Z_{s-}^{(z)},u,y,x)}{Z_{s-}^{(z)}+1}\Big{)}\Big{)}^{2}q({\rm d}u,{\rm d}y,{\rm d}x){\rm d}s\leq\frac{C_{\theta}}{(\ln\theta)^{2}}r,$ (3.60)
for $r\geq 0$. For $p\geq 2$, inequality (3.36) in Lemma 3.9 implies that the (predictable) compensator $(\hat{V}^{(p)}_{r})_{r\geq 0}$ of $(\sum_{s\leq r}|\Delta\hat{Y}_{s}^{(\varepsilon)}|^{p})_{t\geq 0}$ satisfies $\hat{V}^{(p)}_{r}\leq\frac{p!}{2}(\frac{2}{\ln\theta})^{p}C_{\theta}r$, for $r\geq 0$, where $C_{\theta}$ is as in (3.56). Then, it follows from (3.60) and Lemma A.1 that, for $t<T_{D,\varepsilon}(z)\wedge\varepsilon/4\kappa\wedge\eta_{\theta}$,
@@
$\leq\exp\left(-\frac{\varepsilon\ln\theta}{64\sqrt{2}}\right),$ (3.61)
@@
Finally, (3.43) follows by combining (3.3), (3.54), (3.3) and (3.3). ∎
We continue with the proof of Theorem 2.2.
Proof of Theorem 2.2.
@@
$\lim_{t\downarrow 0}\mathbb{E}\left[\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}-1\right|^{p}\right]=0.$ (3.62)
@@
$-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}-1=-\mathbf{m}_{\rm int}s(v_{s}+1)\Big{(}\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}-1\Big{)}-\mathbf{m}_{\rm int}s(v_{s}+1)+s\mathbf{m}_{\rm int}-1$ (3.63)
and thus, (3.62) and Lemma 3.6 imply our claim in Theorem 2.2.
Now we prove (3.62). Note that
$\sup_{s\leq t}\left|\log\left(\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right)\right|=\log\left(\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right|\vee\sup_{s\leq t}\left|\frac{v_{s}+1}{Z_{s}^{(\infty)}+1}\right|\right),\quad t\geq 0.$ (3.64)
Fix $0\leq t<T(\infty)$, and choose $\varepsilon>e$ (i.e., $\ln\varepsilon>1$) such that $t<T(\infty)\wedge(\ln\varepsilon)/4\kappa\wedge\eta_{\theta}$. It follows from Lemma 3.10 that, for $u>\varepsilon$,
$\mathbb{P}\left(\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right|\geq u\right)\leq e^{-\frac{(\ln u)\ln\theta}{64\sqrt{2}}}=u^{-\frac{\log\theta}{64\sqrt{2}}};$ (3.65)
to see this note that $(\ln\varepsilon)/4\kappa\leq(\ln u)/4\kappa$, and thus $t<T(\infty)\wedge(\ln u)/4\kappa\wedge\eta_{\theta}$. By choosing $\theta=\vartheta>e^{128\sqrt{2}}$ as in ($\bf H4$}), we deduce that, for $p\in[1,2]$, there is a constant $0<C_{1}(p)<\infty$ such that
$\mathbb{E}\left[\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}\right|^{p}\right]<C_{1}(p).$ (3.66)
@@
$\mathbb{E}\left[\sup_{s\leq t}\left|\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}-1\right|^{p}\right]<C_{2}(p).$ (3.67)
Therefore, for $p\in[1,2]$, (3.62) follows by Lemma 3.6, (3.34) and the dominated convergence theorem. The case $p\in(0,1)$ follows by Jensen’s inequality. This completes the proof. ∎
Corollary 3.11.
@@
    $\mathbb{E}\Big{[}\sup_{s\leq t}\Big{|}-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}\Big{|}^{p}\Big{]}<C(p),\quad\text{for}\quad t\leq T.$ (3.68)
@@
    $\mathbb{E}\Big{[}\sup_{r\leq s\leq t}Z_{s}^{(\infty)}\Big{]}<\infty\qquad\text{and}\qquad\mathbb{E}\Big{[}\sup_{r\leq s\leq t}(Z_{s}^{(\infty)})^{2}\Big{]}<\infty.$ (3.69)
@@
$-\mathbf{m}_{\rm int}sZ_{s}^{(\infty)}\leq-\mathbf{m}_{\rm int}s(v_{s}+1)\frac{Z_{s}^{(\infty)}+1}{v_{s}+1}.$ (3.70)
Recall that $\mathbf{m}_{\rm int}>0$. By Lemma 3.6, $\lim_{s\downarrow 0}-\mathbf{m}_{\rm int}s(v_{s}+1)=1$. In particular, $s\mapsto-\mathbf{m}_{\rm int}s(v_{s}+1)$ is bounded on $[0,T]$. Then, (3.70), Lemma 3.6 and (3.66) imply (i) for $p\in[1,2]$. For $p\in(0,1)$, the inequality in (i) follows by Jensen’s inequality. To prove (ii), note that, for $0<r\leq s\leq t$,
$Z_{s}^{(\infty)}\leq(-r\mathbf{m}_{\rm int})^{-1}\sup_{r\leq u\leq t}\left|-\mathbf{m}_{\rm int}uZ_{u}^{(\infty)}\right|.$ (3.71)
@@
3.4 Proof of Theorem 2.3
In this section, we prove Theorem 2.3. We follows the approach from [28] (see also [29]), modifying it to account for the branching and cooperative mechanisms of the BPI process. Recall the definition of $T(\infty)$ given before (3.23).
Lemma 3.12.
@@
$Z_{t}^{(\infty)}=Z_{r}^{(\infty)}+\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,x)N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.72)
@@
Recall that ($\bf H4$}) implies $\mathbf{m}_{\rm br}<\infty$, the latter together with $\mathbf{m}_{\rm int}<0$ implies that the BPI process does not explode in finite time; see Example 2.1 in [7]. By Proposition 3.1, for all $z\in\mathbb{N}_{0}$ and $0<r\leq t<T(\infty)$,
@@
$\quad\quad\quad+\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(z)},u,y,x)\mathbf{1}_{\{x\geq 0\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.73)
Proposition 3.2 and Theorem 2.1 imply that, for any $0<t<T(\infty)$, $\lim_{z\rightarrow\infty}Z_{t}^{(z)}=Z_{t}^{(\infty)}$ (increasingly), almost surely. Recall that the Poisson random measure $N$ is an almost-surely $\sigma$-finite measure. Since $z\in\mathbb{R}_{+}\mapsto H(z,u,y,x)\mathbf{1}_{\{x\geq 0\}}$ is an increasing function, the monotone convergence theorem implies that
$\lim_{z\rightarrow\infty}\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(z)},u,y,x)\mathbf{1}_{\{x\geq 0\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)$
$\quad\quad\quad=\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,x)\mathbf{1}_{\{x\geq 0\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.74)
On the other hand, $|H(Z_{s-}^{(z)},u,y,-1)|\mathbf{1}_{\{x=-1\}}\leq|H(Z_{s-}^{(\infty)},u,y,-1)|\mathbf{1}_{\{x=-1\}}$, for all $r\leq s\leq t<T(\infty)$ and $(u,y,x)\in\mathcal{X}$. By Corollary 3.11 (ii), one can easily deduce that $|H(Z_{s-}^{(\infty)},u,y,-1)|\mathbf{1}_{\{x=-1\}}$ is integrable with respect to the Poisson random measure $N$, almost surely. Therefore, by the dominated convergence theorem,
$\lim_{z\rightarrow\infty}\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(z)},u,y,-1)\mathbf{1}_{\{x=-1\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x)$
$\quad\quad\quad=\int_{(r,t]}\int_{\mathcal{X}}H(Z_{s-}^{(\infty)},u,y,-1)\mathbf{1}_{\{x=-1\}}N({\rm d}s,{\rm d}u,{\rm d}y,{\rm d}x).$ (3.75)
@@
Let $\tilde{N}$ be the compensated Poisson point measure $N$ on $\mathbb{R}_{+}\times\mathcal{X}$ defined in (3.3), where $\mathcal{X}\coloneqq\{1,2\}\times(\mathbb{N}\cup\Delta)\times\mathbb{N}_{-1}$.
Lemma 3.13.